OpenStudy (anonymous):

lim x -> 0^+ (1/x)^ xsquare

OpenStudy (anonymous):

\[\lim_{x\to0^+}\left(\frac{1}{x}\right)^{x^2}~~?\]

OpenStudy (anonymous):

yup ..

OpenStudy (anonymous):

l'hospital rule too ~

OpenStudy (jhannybean):

LH rule for powers.. hmm.

OpenStudy (anonymous):

\[\exp\left[\ln\left(\lim_{x\to0^+}\left(\frac{1}{x}\right)^{x^2}\right)\right]\] Continuity of natural log as the argument approaches 0 from the right allows you to "pull out" the limit: \[\exp\left[\lim_{x\to0^+}\ln\left(\frac{1}{x}\right)^{x^2}\right]\] What do you know about logarithms when it comes to exponents like this one?